OpenStudy (anonymous):

What is the solution to the rational equation 4/3 +3/x+5 = 21/3x+15 ? a. x = –8 b. x = –2 c. x = 2 d. x = 8 I keep doing this and then I get stuck... please help!

OpenStudy (anonymous):

\[4/3+3/(x+5)=21/(3x+15)\] Is the question this? Then: \[4/3+3/(x+5)=7/(x+5)\] \[4/3=4/(x+5)\] \[x+5=3\] \[x=-2\] But, is the question that?

OpenStudy (anonymous):

yea i got that too, after now... thanks so b...